OpenStudy (rock_mit182):

Show that (n^2+ n) is divisible by 2 for all natural numbers n.

myininaya (myininaya):

you can factor n^2+n

myininaya (myininaya):

And guess what must be true about one of the factors of n^2+n?

myininaya (myininaya):

or you can also do this by induction

OpenStudy (solomonzelman):

well, yes.... for even n, (n) is even and for odd n (n+1) is even. and Even times Odd = Even (and that means that any output is divisible by 2) (myininaya already posted that)

OpenStudy (rock_mit182):

yeh by induction is what im looking for

OpenStudy (rock_mit182):

could you help me out ?

myininaya (myininaya):

so we are proving this for natural numbers

myininaya (myininaya):

base case is n=1

myininaya (myininaya):

Do you have 2 divides 1^2+1?

myininaya (myininaya):

If so move on to the assumption part of the proof.

myininaya (myininaya):

Suppose that for some integer k>=1 we have 2 divides k^2+k

myininaya (myininaya):

Now the trick is to use that assumption to prove 2 divides (k+1)^2+(k+1)

myininaya (myininaya):

hint: You will need 2a=k^2+k for some integer a. And you will need to multiple out the (k+1)^2 of the (k+1)^2+(k+1) expression .

OpenStudy (rock_mit182):

oh ok let me see

OpenStudy (rock_mit182):

base case : \[k^{2}+k\] now for k+1: \[(k^2+1+(k+1))\]

myininaya (myininaya):

you mean (k+1)^2+(k+1)?

myininaya (myininaya):

base case is proving it for the first case which is n=1

OpenStudy (rock_mit182):

\[k^2 +2k+1+(k+1)\]

myininaya (myininaya):

ok now you are given something about k^2+k

OpenStudy (rock_mit182):

\[k^2+3k+2\] now what Do i have to do .. ?

myininaya (myininaya):

\[k^2+2k+1+(k+1) \\ =k^2+k+2k+1+1=k^2+k+2k+2=k^2+k+2(k+1)\]

myininaya (myininaya):

what are you given about k^2+k

myininaya (myininaya):

Remember you made the assumption that 2a=k^2+k for some integer a.

OpenStudy (rock_mit182):

so 2a +2(k+1) ?

myininaya (myininaya):

yep yep factor the 2 out :)

OpenStudy (rock_mit182):

wow i never thought that my answer would be right lol

myininaya (myininaya):

So you have Base case: n=1: 2 divides 1^2+1 Assumption part: Assume for some integer k>=1 we have 2a=k^2+k for some integer a. Inductive case: (k+1)^2+(k+1)=k^2+k+2(k+1)=2a+2(k+1) =2(a+k+1) which means what?

myininaya (myininaya):

That 2 divides ?

OpenStudy (rock_mit182):

is it beacuase of the coeficient two ?

myininaya (myininaya):

(k+1)^2+(k+1) has the factor 2 and the factor (a+k+1) where a+k+1 is also an integer. having 2 as a factor means that 2 divides the (k+1)^2+(k+1)

OpenStudy (rock_mit182):

wow you're so great at induction

OpenStudy (rock_mit182):

thanks a lot you're amazing

OpenStudy (rock_mit182):

:)

myininaya (myininaya):

The earlier proof was to factor n^2+n and get n(n+1). Knowing n is natural then that means that n or (n+1) will be even since n and (n+1) are consecutive natural numbers. Example 2,3 you have 2 is even 11,10 you have 10 is even so n(n+1) will always have an even factor which means it is divisible by 2

OpenStudy (rock_mit182):

any integer (a +k+1) multyplied by 2 is an even number that's why is divisible by two as well

myininaya (myininaya):

meant to list 10,11 in order oh well 9,10 you have 10 is even

OpenStudy (rock_mit182):

many thanks really really i apreciate your help